|4x - 7| = 1
|3- 8x| = 1
What value of x satisfies both of the equations above?

Answer by stanbon(75887)

|4x - 7| = 1
4x-7 = 1 or 4x-7 = -1
x = 2 or x = 3/2
------------------------

|3- 8x| = 1
3-8x = 1 or 3-8x = -1
8x = 2 or 8x = 4
x = 1/4 or x = 1/2
=====================
Answer to your question: no solution
========================================
Cheers,
Stan H.